OpenStudy (anonymous):

The area of a circle is πr2. To find the area of a sector with a central angle of Θ, measured in radians, by what should you multiply πr2?

OpenStudy (jonnyvonny):

You can set this up as ratio or fraction: area or circle : Degree (in radians) \[(2\pi*r)/2\pi\] The numerator is area of circle, and denominator is the theta in radians. Without the known variables, we may instead put\[A/ \theta\] It doesn't matter whichever is on top, so long as you keep the proper values aligned; 2pi goes with theta and 2pi*r goes with area. \[(2\pi*r)/2\pi = A/ \theta\] From there you may cross multiply and divide. I hope this helped, if it didn't please reply so I may elaborate. :)

OpenStudy (anonymous):

Oh I understand now. Thank you so much!

OpenStudy (anonymous):

I still can't get the final answer! is it theta divided by 2pi?

OpenStudy (jonnyvonny):

Looking at is on computer is confusing, but when you cross multiply it: \[2\pi*r*\theta = 2\pi*A\]

OpenStudy (jonnyvonny):

There are 2 variables in that problem, put to answer your problem, 2pi*r is being multiplied by theta, which is all being divided by 2pi (assuming you desire A, the unknown area, alone).

OpenStudy (anonymous):

Then it is just r * theta = A?